Solve the equation for y. Then find the value for y for each value x.
2y+7x=4; x= 5, 10, 15

Answer by tracyc0912(40)

Plug in each x value and solve for y (you'll get 3 answers!)
2y+7(5)=4
2y+35=4
2y=-31
y=-31/2
2y+7(10)=4
2y+70=4
2y=-66
y=-33
2y+7(15)=4
2y+105=4
2y=-101
y=-101/2
Your 3 answers are (-31/2), (-33), (-101/2)